OpenStudy (anonymous):

Find two functions f(x) and g(x) such that f[g(x)] = x but g[f(x)] does not equal x

OpenStudy (anonymous):

I wouldn't even know where to begin here.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Luigi0210

zepdrix (zepdrix):

I guess we should try to avoid using simple linear functions. A composition of linear functions will always give us g(f(x))=x if it already gave us f(g(x))=x. So we need something a little bit fancier. Like a function that is `not one-to-one`.

OpenStudy (anonymous):

I tried arctan and tan, but she (my teacher) said they'd both come to x. Wolfram alpha says she's wrong though?

zepdrix (zepdrix):

Hmm, ya seems like she's wrong >.< I have another idea.. how bout these two functions :)

zepdrix (zepdrix):

\[\large\rm f(x)=x^2,\qquad\qquad\qquad g(x)=\sqrt{x}\]

OpenStudy (anonymous):

Those both equal x

zepdrix (zepdrix):

\[\large\rm g(f(x))=\sqrt{x^2}=|x|\]

OpenStudy (anonymous):

Ah, nevermind I just saw that

zepdrix (zepdrix):

\[\large\rm f(g(x))=\sqrt{x}^2=x\]When the square is on the outside we have something else going on like we see here :)

zepdrix (zepdrix):

But it's happening because x^2 is not one-to-one. Seems like tangent should give you some similar shenanigans, but i dunno >.<

OpenStudy (anonymous):

Yeah... Thank you! You've helped bunches1

OpenStudy (anonymous):

*!

zepdrix (zepdrix):

cool c: